OpenStudy (anonymous):

Evaluate the expression, using the given value of the variable. 3 – 7x + 2x when x = 3 A. -18 B. -24 C. -30 D. -12

OpenStudy (idealist10):

3-7x+2x=3-5x=3-5(3)=3-15=-12 Answer: -12

OpenStudy (anonymous):

Thank you!